# What Do VLMs NOTICE? A Mechanistic Interpretability Pipeline for Gaussian-Noise-free Text-Image Corruption and Evaluation

### v2

@@ Abstract
Vision-Language Models (VLMs) have gained prominence due to their success in solving complex cross-modal tasks. However, the internal mechanisms of VLMs, particularly the roles of cross-attention and self-attention in multimodal integration, are not fully understood. To address this gap, we introduce NOTICE, a Gaussian-NOise-free Text-Image Corruption and Evaluation pipeline for mechanistic interpretability in VLMs. NOTICE introduces Semantic Image Pairs (SIP) corruption, the first visual counterpart to Symmetric Token Replacement (STR) for text. Through NOTICE, we uncover a set of “universal attention heads” in BLIP and LLaVA that consistently contribute across different tasks and modalities. In BLIP, cross-attention heads implement object detection, object suppression, and outlier suppression, whereas important self-attention heads in LLaVA only perform outlier suppression. Notably, our findings reveal that cross-attention heads perform image-grounding, while self-attention in LLaVA heads do not, highlighting key differences in how VLM architectures handle multimodal learning***Code: https://anonymous.4open.science/r/NOTICE.
@@ 1 Introduction
Figure 1: Universal cross-attention heads in BLIP implement object detection and object suppression.
@@ 2 Related Work
Multimodal Interpretability. Despite the successes of VLMs, the role of attention in multimodal integration remains under-studied Xu et al. (2023). Most works either design probing tasks to identify what vision-language concepts models struggle to capture (Shekhar et al., 2017a; Hendricks and Nematzadeh, 2021; Lu et al., 2023) or assess whether models rely more on one modality over the other Cao et al. (2020); Salin et al. (2022); Hessel and Lee (2020). Probing tasks evaluate whether a model encodes a concept by training a linear classifier on its hidden states. Salin et al. (2022) create visual, text, and multimodal probing tasks to show that models heavily rely on text priors and struggle with object, size, and count concepts. Gradient-based methods analyze gradients to understand how changes in input features influence outputs, assigning importance scores to features and providing insights into the decision-making process. Liang et al. (2023) use second-order gradients, building on LIME Ribeiro et al. (2016) and Shapley values Merrick and Taly (2020), to disentangle cross-modal interactions and determine how much each modality contributes to decisions. While these works have provided insights into multimodal models, recent research shows that 1) probing methods can produce misleading results (Belinkov, 2022) and 2) gradient-based methods for feature attributions can fail to predict behavior (Bilodeau et al., 2024). Instead, Giulianelli et al. (2021); Conneau et al. (2018) demonstrate that causal approaches are more informative than probing representations.
Figure 3: Activation Patching using SIP corruption.
@@ 3 Methods
Image Corruption: Semantic Image Pairs. While STR is a widely used text corruption method, extending semantic corruption to the vision domain presents significant challenges as it requires a semantically aligned image pair. To address this, we propose SIP corruption, a visual corollary to STR, which modifies only one semantic concept in an image (see Figure 2). We implement SIP in three diverse image-classification datasets: SVO-Probes (Hendricks and Nematzadeh, 2021) (an extension to FOIL it! Shekhar et al. (2017b)), MIT States (Isola et al., 2015), and Facial Expressions (Goodfellow et al., 2013). While natural variation in backgrounds can happen between semantic image pairs, we select datasets where nouns, subjects, verbs, and objects are known and can be controlled. The Facial Expressions dataset was curated to include only black-and-white photos of faces expressing different emotions. In MIT-States, object-state pairs are known and can be used to maintain objects (e.g., “building”) while varying their states (e.g., “modern”). The SVO-Probes dataset is designed to vary only one element (subject, verb, or object) at a time (e.g., only varying “dog” and “goat”, while maintaining “grass” and “lying” constant, as seen in Figure 2 panel 1). Thus, for each dataset, we collect the known elements (emotions, objects/states, subjects/verbs/objects) and create image pairs that differ on that element. Our framework can be extended to any image dataset with clearly defined attributes, such as cars, shapes, animals, etc., using the class labels to find clean and corrupt pairs.
@@
Activation Patching. Let $M$ be a vision-language model and $M_{l}$ its hidden states at layer $l\in\{0,1,...,n\}$. The forward pass on an image ($I$) and text ($T$) pair at layer $l$ is denoted as $M_{l}(I,T)$. We define $I^{*}$ and $T^{*}$ as a corrupted image and text. Note that we only corrupt one modality at a time. In a forward pass with $I^{*}$ and $T$, activation patching replaces the hidden states of the corrupted run $M_{l}(I^{*},T)$ with the clean run $M_{l}(I,T)$ to determine which components have the most significant indirect effect on the output. The hidden states at all subsequent layers, $M^{{}^{\prime}}_{j}(I^{*},T)$ for $j\in\{l+1,...,n\}$, are called patched states. In Figure 3, we provide an example illustrating patching using SIP image corruption. The image of the puppy is the clean image, $I$, and the goat is $I^{*}$. Patching the correct answer token “puppy” at $M_{l}(I,T)$, from the clean image to the “puppy” token at $M(I^{*},T)$ creates the patched states $M^{{}^{\prime}}(I^{*},T)$ shown as orange diamonds in Figure 3.
@@ 4 Results
We begin our investigation by conducting module-wise activation patching to understand the global patterns of self-attention, cross-attention, and MLP modules under both image and text corruption. We demonstrate that SIP image corruption produces results closely aligned with STR text corruption and effectively highlights the role of middle-layer MLPs, which have repeatedly proven crucial in a wide range of tasks Meng et al. (2023). Building on this, we perform a fine-grained analysis of cross-attention heads in BLIP and self-attention heads in LLaVA, the critical point of multimodal integration for each model. As we identify the most significant cross-attention heads in BLIP and most significant self-attention heads in LLaVA, we categorize them into three distinct functional roles, offering deeper insights into their contributions and behaviors.
Figure 5: Module-wise activation patching results for SIP, and Gaussian-Noise corruption on SVO-Probes on BLIP.
@@ 4.1 Module-wise Activation Patching
We evaluate the effectiveness of SIP corruption by comparing it with Gaussian-noise, the image corruption technique used in previous work Palit et al. (2023). In addition, we explore semantic image corruption using generative models, which shows the NOTICE framework is generalizable beyond curated datasets like SVO-Probes. Specifically, we use in-painting Stable Diffusion (Rombach et al., 2022) to perform semantic image corruption by generating ’corrupted’ versions of clean images that only vary on one key semantic property. For instance, generating an image of a man sitting on a chair from an image of a man sitting on sand. Figure 5 demonstrates the promise of using generative models to implement SIP corruption since stable-diffusion module-wise patching results are highly similar to SVO-Probes. Note that generative models may occasionally introduce undesirable image artifacts (e.g., adding extra fingers to a hand) that we cannot control. We further analyze this in Appendix D. While using Stable Diffusion to create SIP shows promise for future CMA applications, Gaussian noise proves unreliable. In both LLaVA and BLIP, MLP modules highlight the importance of early and middle layers, while Gaussian noise only highlights the last layer (see Figure 5 and Appendix F.1). Several studies highlight the importance of middle layers in transformers Durrani et al. (2020); Meng et al. (2023), and show that MLPs store syntactic information and task-specific representations often peak here Durrani et al. (2020); Hewitt and Manning (2019); Goldberg (2019); Jawahar et al. (2019). Additionally, self-attention patching in SIP assigns the highest probability to the correct token, unlike Gaussian noise. SIP and STR also display consistent activation patterns—long horizontal regions in MLPs (aligning with Meng et al. (2023)) and vertical regions in self-attention—whereas Gaussian Noise shows clear differences. These findings suggest that SIP is a more effective tool for probing VLMs and identifying key components and interactions.
4.2 Universal Cross-Attention Heads Exist Across Corruption Schemes and Tasks
@@
Model Vision Heads Multimodal Heads Text Heads LLAVA L15.H8 L15.H5, L17.H22 $\emptyset$ L16.H5, L16.H24 L19.H15, L18.H30 BLIP $\emptyset$ L3.H0 L0.H11
Table 1: Universal attention heads in LLaVA and BLIP.
@@
Table 1 shows “universal attention heads”, which we defined as consistently having a logit difference two standard deviations above the average regardless of the task or the modality corruption. In LLaVA, we find no “text only” heads; all important self-attention heads either overlap with multimodal heads that are important for both modalities or emerge under image corruption (“vision only” heads). This suggests that visual inputs still influence the prompt even under text corruption. We hypothesize this is due to LLaVA’s early fusion, where image tokens are embedded into the textual prompt. In contrast, BLIP lacks “vision only” heads, likely because its dual-stream cross-attention design allows the model to develop stronger text-specific learning. These findings highlight key differences in cross-modal integration between the two models.
@@ 4.3 Universal Attention Heads Perform Distinct Functions
To better understand the role of universal attention heads in multimodal integration, we visualize the cross-attention patterns between the ’correct answer’ text token and image patches in BLIP. For LLaVA, we visualize self-attention patterns of universal attention heads from the ’Assistant:’ token to image patches. We find that universal cross-attention heads in BLIP fall into three function classes: object detection, object suppression, and outlier suppression, seen in Figure 1 whereas universal self-attention heads in LLaVA primarily implement outlier suppression seen in Figure  7.
@@
Outlier Suppression. Outlier suppression heads reduce attention to high-probability image patches and distribute attention uniformly across remaining patches. All universal-attention heads in LLaVA implement outlier suppression. In BLIP, L0.H11 implements outlier suppression across all tasks in this paper. For BLIP, L0.H11 is a text-only universal cross-attention head, meaning it does not produce a significant patching effect when images are corrupted. Due to its specialization, L0.H11 does not contribute to the semantic association between text tokens and image patches. Instead, it plays a crucial role in filtering out irrelevant visual information, enhancing the model’s robustness.

### v3

@@ Abstract
Vision-Language Models (VLMs) have gained prominence due to their success in solving complex cross-modal tasks. However, the internal mechanisms of VLMs, particularly the roles of cross-attention and self-attention in multimodal integration, are not fully understood. To address this gap, we introduce NOTICE, a Gaussian-NOise-free Text-Image Corruption and Evaluation pipeline for mechanistic interpretability in VLMs. NOTICE introduces Semantic Image Pairs (SIP) corruption, the first visual counterpart to Symmetric Token Replacement (STR) for text. Through NOTICE, we uncover a set of “universal attention heads” in BLIP and LLaVA that consistently contribute across different tasks and modalities. In BLIP, cross-attention heads implement object detection, object suppression, and outlier suppression, whereas important self-attention heads in LLaVA only perform outlier suppression. Notably, our findings reveal that cross-attention heads perform image-grounding, while self-attention in LLaVA heads do not, highlighting key differences in how VLM architectures handle multimodal learning***Code: https://github.com/wrudman/NOTICE/.
@@ 2 Related Work
Multimodal Interpretability. Despite the successes of VLMs, the role of attention in multimodal integration remains under-studied Xu et al. (2023). Most works either design probing tasks to identify what vision-language concepts models struggle to capture (Shekhar et al., 2017a; Hendricks and Nematzadeh, 2021; Lu et al., 2023) or assess whether models rely more on one modality over the other Cao et al. (2020); Salin et al. (2022); Hessel and Lee (2020). Probing tasks evaluate whether a model encodes a concept by training a linear classifier on its hidden states. Salin et al. (2022) create visual, text, and multimodal probing tasks to show that models heavily rely on text priors and struggle with object, size, and counting. Gradient-based methods analyze gradients to understand how changes in input features influence outputs, assigning importance scores to features and providing insights into the decision-making process. Liang et al. (2023) use second-order gradients, building on LIME Ribeiro et al. (2016) and Shapley values Merrick and Taly (2020), to disentangle cross-modal interactions and determine how much each modality contributes to decisions. While these works have provided insights into multimodal models, recent research shows that 1) probing methods can produce misleading results (Belinkov, 2022) and 2) gradient-based methods for feature attributions can fail to predict behavior (Bilodeau et al., 2024). Instead, Giulianelli et al. (2021); Conneau et al. (2018) demonstrate that causal approaches are more informative than probing representations.
Figure 2: Activation Patching using SIP corruption. The image of the puppy is the clean image, $I$, and the goat is the corrupt image, $I^{*}$. Patching the correct answer token “puppy” at $M_{l}(I,T)$, from the clean image to the “puppy” token at $M(I^{*},T)$ creates the patched states $M^{{}^{\prime}}(I^{*},T)$ shown as orange diamonds.
@@ 3 Methods
Model Selection For our model selection, we use LLaVA-1.5-7B (Vicuna base) Liu et al. (2024) and BLIP-VQA-Base Li et al. (2022), chosen for their relevance and alignment with recent interpretability research. BLIP-VQA was included to enable direct comparison with prior work applying activation patching to vision-language models Palit et al. (2023), facilitating a meaningful evaluation of our corruption scheme (NOTICE) against Gaussian-noise patching. In addition, we also focus on LLaVA-1.5 to align with more recent interpretability studies, including Luo et al. (2024), Neo et al. (2024), and Jiang et al. (2024).
@@
Image Corruption: Semantic Image Pairs. While STR is a widely used text corruption method, extending semantic corruption to the vision domain presents significant challenges as it requires a semantically aligned image pair. To address this, we propose SIP corruption, a visual corollary to STR, which modifies only one semantic concept in an image (see Figure 1). We implement SIP in three diverse image-classification datasets: SVO-Probes (Hendricks and Nematzadeh, 2021) (an extension to FOIL it! Shekhar et al. (2017b)), MIT States (Isola et al., 2015), and Facial Expressions (Goodfellow et al., 2013). While natural variation in backgrounds can happen between semantic image pairs, we select datasets where nouns, subjects, verbs, and objects are known and can be controlled. The Facial Expressions dataset was curated to include only black-and-white photos of faces expressing different emotions. In MIT-States, object-state pairs are known and can be used to maintain objects (e.g., “building”) while varying their states (e.g., “modern”). The SVO-Probes dataset is designed to vary only one element (subject, verb, or object) at a time (e.g., only varying “puppy” and “goat”, while maintaining “grass” and “lying” constant, as seen in Figure 1 panel 1). Thus, for each dataset, we collect the known elements (emotions, objects/states, subjects/verbs/objects) and create image pairs that differ on that element. Our framework can be extended to any image dataset with clearly defined attributes, such as cars, shapes, animals, etc., using the class labels to find clean and corrupt pairs.
@@
Activation Patching. Let $M$ be a vision-language model and $M_{l}$ its hidden states at layer $l\in\{0,1,...,n\}$. The forward pass on an image ($I$) and text ($T$) pair at layer $l$ is denoted as $M_{l}(I,T)$. We define $I^{*}$ as a corrupted image and $T^{*}$ as a corrupted text input. Note that we only corrupt one modality at a time. In a forward pass with $I^{*}$ and $T$, activation patching replaces the hidden states of the corrupted run $M_{l}(I^{*},T)$ with the clean run $M_{l}(I,T)$ to determine which components have the most significant indirect effect on the output. The hidden states at all subsequent layers, $M^{{}^{\prime}}_{j}(I^{*},T)$ for $j\in\{l+1,...,n\}$, are called patched states. In Figure 2, we provide an example illustrating patching using SIP image corruption. The image of the puppy is the clean image, $I$, and the goat is $I^{*}$. Patching the correct answer token “puppy” at $M_{l}(I,T)$, from the clean image to the “puppy” token at $M(I^{*},T)$ creates the patched states $M^{{}^{\prime}}(I^{*},T)$ shown as orange diamonds in Figure 2.
Figure 4: Module-wise activation patching results for SIP, and Gaussian-Noise corruption on SVO-Probes on BLIP. SIP corruption produces activation patterns that align with Stable Diffusion results and highlight the importance of middle layers, while Gaussian noise fails to reveal meaningful attention layers, emphasizing the effectiveness of SIP for probing vision-language models.
@@ 4 Results
We begin our investigation by conducting module-wise activation patching to understand the global patterns of self-attention, cross-attention, and MLP modules under both image and text corruption. We demonstrate that SIP image corruption produces results closely aligned with STR text corruption and effectively highlights the role of middle-layer MLPs, which have repeatedly proven crucial in a wide range of tasks Meng et al. (2023). In contrast, Gaussian Noise image corruption only identifies the final MLP layers as producing a significant patching effect. Building on this, we use SIP to perform a fine-grained analysis of cross-attention heads in BLIP and self-attention heads in LLaVA, the critical point of multimodal integration for each model. As we identify the most significant cross-attention heads in BLIP and most significant self-attention heads in LLaVA, we categorize them into three distinct functional roles, offering deeper insights into their contributions and behaviors.
@@ 4.1 Module-wise Activation Patching
We evaluate the effectiveness of SIP corruption by comparing it with Gaussian-noise, the image corruption technique used in previous work Palit et al. (2023). In addition, we explore semantic image corruption using generative models, which shows the NOTICE framework is generalizable beyond curated datasets like SVO-Probes. Specifically, we use in-painting Stable Diffusion (Rombach et al., 2022) to perform semantic image corruption by generating “corrupted” versions of clean images that only vary on one key semantic property. For instance, generating an image of a man sitting on a chair from an image of a man sitting on sand. Figure 4 demonstrates the promise of using generative models to implement SIP corruption since stable-diffusion module-wise patching results are highly similar to SVO-Probes. Note that generative models may occasionally introduce undesirable image artifacts (e.g., adding extra fingers to a hand) that we cannot control. We further analyze this in Appendix D. While using Stable Diffusion to create SIP shows promise for future CMA applications, Gaussian noise proves unreliable. In both LLaVA and BLIP, MLP modules highlight the importance of early and middle layers, while Gaussian noise only highlights the last layer (see Figure 4 and Appendix F.1). Several studies highlight the importance of middle layers in transformers Durrani et al. (2020); Meng et al. (2023), and show that MLPs store syntactic information and task-specific representations often peak in middle layer MLPs Durrani et al. (2020); Hewitt and Manning (2019); Goldberg (2019); Jawahar et al. (2019). Additionally, self-attention patching in SIP assigns the highest probability to the correct token, unlike Gaussian noise. SIP and STR also display consistent activation patterns —long horizontal regions in MLPs (aligning with Meng et al. (2023)) and vertical regions in self-attention —whereas Gaussian Noise shows clear differences. These findings suggest that SIP is a more effective tool for probing VLMs and identifying key components and interactions.
@@
4.2 Universal Attention Heads Exist Across Corruption Schemes and Tasks
Model Vision Heads Multimodal Heads Text Heads LLaVA L15.H5, L16.H18 L28.H7, L31.H27 L22.H16, L29.H19 L14.H4, L15.H8 L16.H24, L16.H24 L18.H26, L17.H0 L19.H15, L18.H30 L17.H22, L14.H28 L18.10, L21.H30 L19.H4, L15.H14 L17.H13, L20.H28 BLIP L5.H3 L3.H0 L0.H11
Table 1: Universal attention heads in LLaVA and BLIP, identified as having consistently high logit differences across tasks and modality corruptions. In LLaVA, most critical self-attention heads serve multimodal or vision-specific roles. In contrast, BLIP’s dual-stream design results in distinct vision, text, and multimodal heads, supporting more modular cross-modal integration.
@@ 4.2 Universal Attention Heads Exist Across Corruption Schemes and Tasks
Table 1 shows “universal attention heads”, which we defined as consistently having a logit difference two standard deviations above the average regardless of the task or the modality corruption. In LLaVA, we find only find two“text only” heads; a majority of important self-attention heads either overlap with multimodal heads that are important for both modalities or emerge under image corruption (“vision only” heads). This suggests visual inputs still influence the prompt, even under text corruption. We hypothesize this is due to LLaVA’s early fusion, where image tokens are embedded into the textual prompt. In contrast, BLIP has exactly one vision, one text, and one multimodal head, likely because its dual-stream cross-attention design allows the model to fuse visual and textual information more effectively without requiring multiple specialized heads for different modalities. These findings highlight key differences in cross-modal integration between the two models.
Figure 6: Universal cross-attention heads in BLIP implement object detection and object suppression. We visualize the average of all attention heads compared to the universal heads performing each function.
@@ 4.3 Universal Attention Heads Perform Distinct Functions
To better understand the role of universal attention heads in multimodal integration, we visualize the cross-attention patterns between the “correct answer” text token and image patches in BLIP. For LLaVA, we visualize self-attention patterns of universal multimodal self-attention heads from the “Assistant:” token to image patches. We find that universal cross-attention heads in BLIP fall into three function classes: object detection, object suppression, and outlier suppression, seen in Figure 6 whereas universal multimodal self-attention heads in LLaVA primarily implement outlier suppression seen in Figure  7.
@@
Outlier Suppression. Outlier suppression heads reduce attention to high-probability image patches and distribute attention uniformly across remaining patches. All universal multimodal attention heads in LLaVA implement outlier suppression. In BLIP, L0.H11 implements outlier suppression across all tasks in this paper. For BLIP, L0.H11 is a text-only universal cross-attention head, meaning it does not produce a significant patching effect when images are corrupted. Due to its specialization, L0.H11 does not contribute to the semantic association between text tokens and image patches. Instead, it plays a crucial role in filtering out irrelevant visual information, enhancing the model’s robustness.